# Disentangled Representation with Cross Experts Covariance Loss for Multi-Domain Recommendation

### v1

@@ Abstract.
Multi-domain learning (MDL) has emerged as a prominent research area aimed at enhancing the quality of personalized services. The key challenge in MDL lies in striking a balance between learning commonalities across domains while preserving the distinct characteristics of each domain. However, this gives rise to a challenging dilemma. On one hand, a model needs to leverage domain-specific modules, such as experts or embeddings, to preserve the uniqueness of each domain. On the other hand, due to the long-tailed distributions observed in real-world domains, some tail domains may lack sufficient samples to fully learn their corresponding modules. Unfortunately, existing approaches have not adequately addressed this dilemma. To address this issue, we propose a novel model called Crocodile, which stands for Cross-experts Covariance Loss for Disentangled Learning. Crocodile adopts a multi-embedding paradigm to facilitate model learning and employs a Covariance Loss on these embeddings to disentangle them. This disentanglement enables the model to capture diverse user interests across domains effectively. Additionally, we introduce a novel gating mechanism to further enhance the capabilities of Crocodile. Through empirical analysis, we demonstrate that our proposed method successfully resolves these two challenges and outperforms all state-of-the-art methods on publicly available datasets. We firmly believe that the analytical perspectives and design concept of disentanglement presented in our work can pave the way for future research in the field of MDL.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Crocodile: Cross Experts Covariance for Disentangled Learning in Multi-Domain Recommendation; October 21-25, 2024; Boise, Idaho, USAisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
((a)) Dillema of Interest Diversity v.s. Collapse Prevention
@@
Figure 1. The proposed Crocodile successfully resolves the dilemma of interest diversity v.s. collapse prevention and achieves the best performance regarding gAUC.
@@
Online recommendation systems (Broder, 2008) are now essential tools for businesses to effectively reach their target audiences, servicing billions of internet users daily. Empowered by advanced algorithms and vast traffic, these systems play a crucial role in modeling user experiences and generating revenue for advertisers. However, traditional methods still face substantial challenges due to the dynamic nature of online platforms, which are marked by diverse user preferences across domains. To address these challenges, researchers have increasingly turned to multi-domain learning (MDL) (Chang et al., 2023; Zhou et al., 2023; Sheng et al., 2021; Yang et al., 2022; Li et al., 2023) as a promising approach to enhance the performance by involving information from multiple related domains.
Unfortunately, MDL for recommendation often confronts the challenge of preference conflict, where a user might exhibit divergent behaviors across domains. To remedy this, the MDL community aims to develop domain-specific components, including STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and PEPNet (Chang et al., 2023) have progressively integrated domain-specific components.
Interestingly, we observe a trend in MDL methods towards the development of bottom-level domain-specific modules. In this work, we uncover the mechanisms behind this trend. To the best of our knowledge, this is the first study to reveal the mechanism of increasing interest diversity to resolve preference conflict. Furthermore, we propose a novel metric, the Diversity Index (DI), to quantify interest diversity, as shown in Fig 1(a).
Based on our discovery, we notice that domain-specific embedding (Su et al., 2023; Guo et al., 2023a) incurs the highest diversity and hence should be able to resolve preference conflict. However, naively deploying domain-specific embeddings does not turn out to be fruitful, we observe negative transfer in domains with fewer data. Delving further, we find out the core issue is sub-optimal embeddings for the domain with less data, i.e., they are dimensionally collapsed. [^1] We adopt the note of dimensional collapse in the work of Chi et al. (2022) and utilize information abundance (IA) (Guo et al., 2023b) to quantify dimensional collapse.
As shown in Fig. 1(a), our analysis demonstrates that existing works struggle to maintain these two properties: maintaining diversity and avoiding dimensional collapse, thus resulting in the sub-optimal performance shown in Fig. 1(b). As a result, it remains an open question whether we can incorporate both properties.
To this end, we propose a novel a novel Cross-experts Covariance Loss for Disentangled Learning model (Crocodile) framework, which we build this Covariance Loss (CovLoss) to alleviate embedding collapse while in the meantime enabling learning diverse user interests across various domains. To further unleash the power of Crocodile, we develop a novel Prior Informed element-wise Gating (PeG) mechanism to work with CovLoss.
@@
Empirical analysis on two public datasets confirms that Crocodile achieves state-of-the-art results. Further analysis also demonstrates that Crocodile can simultaneously mitigate embedding collapse and learn diverse representations. Ablation experiments reveal that each proposed component contributes positively to the overall performance. Our contributions can be briefly summarized as follows:
  • We discover the mechanism of previous works in maintaining diversity, especially the trend of focusing more on bottom-level modules. We propose a novel metric, Diversity Index (DI), to quantify the diversity and conclude that the trend of previous works is optimizing to increase DI.
  • We reveal that existing works fall into the dilemma of maintaining diversity and avoiding collapse, hence resulting in inferior performance.
@@
  • We propose the Crocodile approach, which successfully balances both properties. Empirical evidence on two public datasets demonstrates that Crocodile achieves state-of-the-art performance and resolves the aforementioned dilemma through comprehensive analysis.
2. Related Work
2.1. Multi-domain Recommendation
Multi-domain learning (MDL) is a promising topic that has garnered attention from industrial recommendations. Many MDL methods have made great efforts to model complex multi-domain interests. Among them, MoE-based methods are widely applied in industrial recommendation systems, such as Taobao (Sheng et al., 2021) , Meituan (Chang et al., 2023) and Kuaishou (Zhou et al., 2023).
STAR (Sheng et al., 2021) utilizes the dot product of domain-specific and domain-shared weight matrices to capture domain-specific interests. HiNet (Zhou et al., 2023) proposes two expert-like components to capture cross-domain correlations and within-domain information explicitly. PEPNet (Chang et al., 2023) introduces PPNet and EPNet, which incorporate domain information through gating. These methods often innovate from the perspective of making more bottom-level modules be domain-personalized within the model.
Recently, STEM (Su et al., 2023) proposed a multi-embedding (ME) structure in multi-task learning (MTL) and step forward in allocating more bottom-level domain-personalized structure, i.e., embedding. It achieved better results compared to both single and multi-embedding versions of MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) in MTL experiments. However, STEM fails in the multi-domain recommendation. We will discuss the reasons in detail in the subsequent sections.
Besides, some other paradigms are applied for multi-domain recommendation. AdaSparse (Yang et al., 2022) and ADL (Li et al., 2023) applied adaptive learning mechanism. EDDA (Ning et al., 2023a), COAST (Zhao et al., 2023) and UniCDR(Cao et al., 2023) provided graph learning-based solutions.
2.2. Disentangled Representation Learning
In the realm of disentangled representation learning (DRL), methods based on spatial or statistical relationships have been proposed.
On the one hand, methods based on spatial relationships achieve disentangling by manipulating the distances or angles between representations. OLE (Lezama et al., 2018), dot product(Wu et al., 2021) and cosine similarity(Liu et al., 2021) push representations towards orthogonal. GAN-based models (Chen et al., 2016; Yu and Welch, 2021; Khrulkov et al., 2021) promote the discrimination among representations. However, they are computationally expensive and not conducive to online recommendation.
On the other hand, statistical methods view representations as instances drawn from high-dimensional distributions and design loss based on statistical measures. Distance correlation (dCorr)(Liu et al., 2022a) encourages lower dCorr between different factors. ETL (Chen et al., 2023) models the joint distribution of user behaviors across domains with equivalent transformation. VAE-based methods (Ma et al., 2019; Wang et al., 2022; Liu et al., 2022b) encourage higher KL divergence between dimensions. However, they often assume a normal prior distribution, which is challenging to satisfy in the context of sparse interactions (Rendle, 2010), especially under sample volume imbalance of multi-domain recommendation.
Besides, some methods integrate both aspects mentioned above. VICReg (Bardes et al., 2021; Shwartz-Ziv et al., 2024) employs covariance and variance as loss components within representations to promote independence across dimensions and prevent collapse while utilizing Euclidean distance between representations to ensure semantic similarity among representations of different patches of an image.
2.3. Dimensional Collapse
Dimensional collapse(Chi et al., 2022) is defined as representations that end up spanning a lower-dimensional space, which is evident through the pronounced imbalance in the distribution of singular values. Chi et al. (2022) found that SMoE structure (Fedus et al., 2022; Lepikhin et al., 2020) may lead to dimensional collapse due to the influence of gating. The work of (Guo et al., 2023b) claimed that this problem is also observed in popular models (e.g., DCN V2 (Wang et al., 2021) and Final MLP (Mao et al., 2023)). To measure the collapse, they proposed the Information Abundance (IA) (Guo et al., 2023b) metric, defined as the ratio between the sum and the maximum l1-norm value of the target singular values. A small IA indicates severe insignificance of tail dimensions and potential dimensional collapse.
Beyond a single perspective, some methods (Hua et al., 2021) built the bridge between representation disentanglement and dimensional collapse prevention. It also claimed feature decorrelation helps alleviate dimensional collapse by reducing strong correlations between axes and promoting more diverse and informative representations.
@@ 3. On the Challenges of Multi-Domain Recommendation
As we have discussed in Sec. 1, existing works fall into the dilemma of avoiding dimensional collapse and maintaining interest diversity. In this section, we will discuss these two properties thoroughly.
3.1. Preventing Dimensional Collapse
According to Fig 1(a), domain-specific embeddings demonstrate its effectiveness in capturing diverse user interests. However, the naive adoption of domain-specific embeddings does not yield the expected outcomes. In fact, the methods with domain-specific embeddings, i.e., ME-PLE (Tang et al., 2020; Su et al., 2023) and STEM (Su et al., 2023), both performance deteriorate in the domain with less data. For example, in domain S6 of Kuairand1k (Gao et al., 2022) dataset, the ME-PLE and STEM show -0.4% and -0.1% lower AUC than the single domain method. This phenomenon has sparked our deep investigation into its underlying mechanism.
Intuitively, the domain-specific embeddings are trained solely on data from the corresponding domain, which results in insufficient samples for domains with less data. Insufficient training data fails to provide adequate variance for high-dimensional representations, resulting in high-dimensional representations residing within a lower-dimensional subspace, called dimensional collapse (Guo et al., 2023b; Chi et al., 2022; Hua et al., 2021).
To measure the collapse, we adopt the IA metric (Guo et al., 2023b), which is defined as $IA(E)=||\sigma||_{1}/||\sigma||_{\infty}$, where $E$ is the embedding of feature (e.g., item id) and $||\cdot||_{1},||\cdot||_{\infty}$ represent the $l1$-norm and $\infty$-norm, respectively.
Remark 1.
Intuitively, singular values $\sigma_{i}$ represent the significance of $i$-th singular value, and IA captures the balance of the singular value distribution. Thus, a small IA indicates severe insignificance of tail dimensions and potential dimension collapse.
@@ 3.1. Preventing Dimensional Collapse
Figure 2. Information Abundance ($log(IA)$) dynamics of STEM’s bottom layer of experts and item embeddings on Kuairand1k. High and Low denote the items with the highest and lowest frequencies, respectively.
Taking STEM as the example of methods applying domain-specific embedding, we analyzed its training IA dynamics of experts and embeddings in Fig. 2. In Kuairand1k, S1 and S6 are the domains with the most and least data. In Fig. 2(a), the S6-specific expert exhibits a lower IA compared to S1, indicating that domains with a smaller data volume are more susceptible to dimensional collapse at the expert level. In Fig. 2(b), we further categorized items into five groups based on their exposure frequency and observed the collapse of item ID embeddings for both the highest and lowest frequency groups. We discovered that domain S6 exhibits a lower IA than the overall S6, indicating a more severe collapse, regardless of the group. Within each domain, the low-exposure group consistently shows a lower IA than the high-exposure group. This further corroborates the notion that a lower exposure count (i.e., data volume) is more prone to dimensional collapse.
Both analytical and empirical evidence suggest that inadequate data of domain can lead to dimensional collapse of the domain-specific embedding, hence leading to inferior performance.
3.2. Interest Diversity
Intuitively, the ability to capture various user interests across domains can be directly characterized by the diversity of representations. When facing conflict preference across domains, a set of diverse representations is enabled to encode both prefer/non-prefer semantics and thus provide distinct vectors to towers of those conflict domains.
To this end, we propose a novel metric, the Diversity Index (DI), to quantify the representations’ ability to capture conflict preference across domains. Specifically, DI depicts the ratio of samples that contain both large and small norm of representations within the whole conflict samples, which is formulated as [^2] Here $AF$ and $SF$ are abbreviations for all forward and specific forward respectively, aligned with the concepts in  (Su et al., 2023).:
$\mathcal{D}_{con(i,j)}$ contains preference conflict samples between domain $i$ and $j$. $\mathbb{I}[\text{{condition}}]$ is the indication function, which is assigned to 1 if the condition is true; otherwise, 0. $M$ represents experts, $\mathbf{O}_{k}^{(p)}$ means the output of $p$-th expert and $k$-th sample, $\tau_{t}$ and $\tau_{b}$ are the thresholds of big and small l2-norm.
Remark 2.
As shown in Fig. 1(a), ME-MMoE (Guo et al., 2023a) exhibits significantly lower diversity compared to STEM (Su et al., 2023), and its improvement in diversity is relatively limited when compared to MMoE (Ma et al., 2018b). This suggests that the representations learned by ME-MMoE (Guo et al., 2023a) lack sufficient ability to solve preference conflict challenges.
@@ 3.2. Interest Diversity
4. Methodology
We propose a multi-embedding architecture with Cross-experts Covariance Loss for Disentangled Learning (Crocodile) to resolve the dilemma, containing Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG), as shown in Fig. 3.

### v2

# Crocodile: Cross Experts Covariance for Disentangled Learning in Multi-Domain Recommendation

@@ Abstract.
Multi-domain learning (MDL) has become a prominent topic in enhancing the quality of personalized services. It’s critical to learn commonalities between domains and preserve the distinct characteristics of each domain. However, this leads to a challenging dilemma in MDL. On the one hand, a model needs to leverage domain-aware modules such as experts or embeddings to preserve each domain’s distinctiveness. On the other hand, real-world datasets often exhibit long-tailed distributions across domains, where some domains may lack sufficient samples to effectively train their specific modules. Unfortunately, nearly all existing work falls short of resolving this dilemma. To this end, we propose a novel Cross-experts Covariance Loss for Disentangled Learning model (Crocodile), which employs multiple embedding tables to make the model domain-aware at the embeddings which consist most parameters in the model, and a covariance loss upon these embeddings to disentangle them, enabling the model to capture diverse user interests among domains. Empirical analysis demonstrates that our method successfully addresses both challenges and outperforms all state-of-the-art methods on public datasets. During online A/B testing in Tencent’s advertising platform, Crocodile achieves 0.72% CTR lift and 0.73% GMV lift on a primary advertising scenario.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
((a)) Dillema of Preserving Distinctiveness v.s. Sufficient Parameters Learning
@@
Figure 1. Our proposed Crocodile successfully resolves the dilemma of preserving domain distinctiveness (measured by Diversity Index) v.s. sufficient parameters learning (measured by Information Abundance), and achieves the best performance measured by gAUC.
@@
Online recommendation systems (Broder, 2008) are now essential tools for businesses to effectively reach their target audiences, servicing billions of internet users daily. Empowered by advanced algorithms and vast traffic, these systems play a crucial role in modeling user experiences and generating revenue for advertisers. However, traditional methods still face substantial challenges due to the dynamic nature of online platforms, which are marked by diverse user preferences across domains. To address these challenges, researchers have increasingly turned to multi-domain learning (MDL) (Chang et al., 2023; Zhou et al., 2023; Sheng et al., 2021; Yang et al., 2022; Li et al., 2023b) as a promising approach to enhance the performance by involving information from multiple related domains.
However, MDL faces conflicts of interest among users across different domains, which limits the potential for further exploitation of multi-domain information. Consequently, existing MDL models attempt to maintain domain distinctiveness through the implementation of domain-aware components along the upside-down trend, i.e., from Tower to Embedding. Specifically, Shared-Bottom (Caruana, 1997) and STAR (Sheng et al., 2021) focuses on making the tower domain-specific, while HiNet (Zhou et al., 2023), M-scan (Zhu et al., 2024) and HierRec (Gao et al., 2024) shift the focus to the expert level. PEPNet (Chang et al., 2023) follows this trend further by emphasizing domain-specific transformed embeddings. However, due to the domination of embeddings regarding the number of parameters in recommendation (Pan et al., 2024), these experts or transformation networks have limited capacities, making their ability to preserve each domain’s distinctiveness in doubt (refer to Finding 1 in Sec. 2.1).
Meanwhile, recent work in recommendation shows the efficacy of learning multiple separate embeddings for each feature. In particular, the Multi-Embedding paradigm (Guo et al., 2023a; Pan et al., 2024) was proposed in single-task learning for feature interaction models, aiming to resolve the dimensional collapse of embeddings caused by the Interaction-Collapse theory (Guo et al., 2023a). Regarding multi-task learning, a variant of multiple embeddings, namely the Shared and Task-specific EMbedding (STEM) (Su et al., 2023) was proposed to retain each task’s distinct characteristics via task-specific embeddings.
Following these works, we try incorporating Shared and Domain-specific EMbeddings (SDEM) in multi-domain learning. However, these models are still not performing well. After careful analysis, we found that the embeddings are insufficiently learned, especially for those small domains (refer to Finding 2 in Sec. 2.2). These small domains have much fewer training samples, making it challenging to learn independent embedding tables for them.
In this paper, we propose a novel Diverse Multi-Embeddings paradigm, which adopts multiple shared embeddings with a novel disentanglement loss. Every embedding table is shared and hence optimized by all domains, overcoming the above-mentioned challenge of embedding insufficient learning, while the disentanglement loss promotes the diversity of these embeddings to preserve the underlying characteristics of all domains, as shown in Fig. 1(a). Upon these multiple disentangled embeddings, a novel Prior Informed element-wise Gating (PeG) mechanism was proposed to route between these embeddings and domain-specific classification towers. We name our method Crocodile, standing for Cross-experts Covariance loss for Disentangled Learning.
@@
Empirical analysis conducted on two public datasets demonstrates that Crocodile achieves state-of-the-art results, as briefly shown in Fig. 1(b). Further analysis reveals that Crocodile effectively mitigates embedding collapse while simultaneously learning diverse representations. Ablation experiments confirm the positive contributions of each proposed component to the overall performance. Online deployment on Tencent’s advertising platform shows a significant CTR and GMV lift. In summary, our contributions can be summarized as follows:
  • We discover a dilemma between learning domain-specific characteristics and sufficient learning of parameters of existing multi-domain learning methods.
  • We propose the Crocodile approach, which employs a Disentangled Multi-Embedding paradigm and a Prior Informed element-wise Gating (PeG) mechanism.
@@
  • Empirical evidence on two public datasets as well as online A/B testing on Tencent’s advertising platform demonstrate that Crocodile achieves state-of-the-art performance and resolves the aforementioned dilemma.
@@ 2. On the Challenges of Multi-Domain Recommendation
In this section, we study to what extent existing multi-domain recommendation models resolve the following two challenges: preserving each domain’s distinctiveness and sufficient learning of embeddings. Besides existing works, inspired by the success of multi-embedding (Su et al., 2023; Guo et al., 2023a), we developed two new baselines: a Shared and Domain-specific EMbeddings (SDEM) approach, which corresponds to the STEM in MTL, a Multi-Embedding PLE for multi-domain learning, ME-PLE in short, which corresponds to the ME-PLE in multi-task learning, and a Multi-Embedding MMoE for multi-domain learning, ME-MMoE in short. In the following, we evaluate the ability of existing and proposed methods to preserve each domain’s distinctiveness and the learning sufficiency of parameters.
2.1. Preserving Distinctiveness
Intuitively, the ability to capture various user interests across domains can be directly characterized by the diversity of representations. When facing conflict preference across domains, a set of diverse representations is enabled to encode both prefer/non-prefer semantics and thus provide distinct vectors to towers of those conflict domains. To this end, we propose a novel Diversity Index (DI) metric to quantify the diversity among the outputs of experts.
We select a subset of samples where there is a diverse interest across domains, quantified by the output norm of domain-specific experts. We denote the set of samples satisfying such conditions as $\mathcal{D}_{(i,j)}$. Formally,
(1) $\mathcal{D}_{(i,j)}=\{k|\|\bm{O}\|^{i}_{\text{SD}}(x_{k})\geq\tau_{t}\cap\|\bm{O}\|^{j}_{\text{SD}}(x_{k})\leq\tau_{b}\},$
where $\tau_{t}$ and $\tau_{b}$ are two specified norm thresholds.
Then, for those models where no experts are bound to any domains such as MMoE or STAR, we simply define the Diversity Index by whether there is one expert that learns a strong interest whilst another one learns a weak interest. That is, we just need to count the ratio of samples that have both a strong expert output norm and a weak one:
(2) $\frac{\sum\limits_{k\in\mathcal{D}_{(i,j)}}\mathbb{I}[\exists p,q,(\|\mathbf{O}^{p}(x_{k})\|_{2}\geq\tau_{t})\cap(\|\mathbf{O}^{q}(x_{k})\|_{2}\leq\tau_{b})]}{|\mathcal{D}_{(i,j)}|},$
where $\mathbb{I}[\text{{condition}}]$ is the indication function, which is assigned to 1 if the condition is true; otherwise, 0. $p,q$ denote the index of experts, $\mathbf{O}^{p}(x_{k})$ means the output of $p$-th expert of sample $x_{k}$.
For those models with domain-specific experts, such as PLE, SDEM, and ME-PLE, we can quantify the Diversity Index by counting the ratio of samples that have a large expert output norm on domain $i$’s specific expert while a small norm on $j$’s expert:
(3) $\frac{\sum\limits_{k\in\mathcal{D}_{(i,j)}}\mathbb{I}[(\|\mathbf{O}^{i}(x_{k})\|_{2}\geq\tau_{t})\cap(\|\mathbf{O}^{j}(x_{k})\|_{2}\leq\tau_{b})]}{|\mathcal{D}_{(i,j)}|}.$
After we compute the diversity measurement $DI_{(i,j)}$ for all domain pairs $(i,j)$, we then take an average of them, getting the overall Diversity Index: $DI=\text{avg}(DI_{(i,j)})$. We then calculate the diversity index of several existing multi-domain recommendation models and those proposed models with multiple embeddings, as illustrated in Fig. 1(a). We have the following findings.
First, existing methods, in general, learn similar interests across domains, as indicated by the low diversity index. For example, the diversity index of HiNET, STAR, PEPNet, and PLE are less than 0.3. That means, for those samples that have shown strong domain-wise interest distinctiveness, these models can capture such distinctiveness for less than 30% of them. MMoE performs the best, with a diversity index of 0.615, but it’s still not a satisfactory result. Second, those multi-embedding models have higher diversity indices, especially compared to their single-embedding correspondence. ME-PLE gets a diversity index 0.403, compared to 0.284 of its single-embedding variant PLE; ME-MMoE gets a diversity index 0.716, compared to 0.615 of its single-embedding variant MMoE. SDEM gets the highest diversity index among models, i.e., 0.904. We conclude with the following finding:
Finding 1. Existing single-embedding recommendation models for multi-domain learning fail to preserve domain distinctiveness. When employing multiple embeddings, especially domain-specific embeddings, models can preserve domain distinctiveness more decently.
2.2. Sufficient Parameter Learning
We found that the proposed ME-PLE and SDEM models perform poorly, particularly in domains with less data. For instance, in domain S6 of the Kuairand1k dataset (Gao et al., 2022), ME-PLE and SDEM show -0.4% and -0.1% lower AUC than the single domain method.
We investigated the learned parameters, focusing on the embeddings. we assess the quality of embedding learning, particularly focusing on dimensional collapse testing, by analyzing the distribution of singular values and employing the aggregation metric Information Abundance (IA) (Guo et al., 2023b). It is defined as the sum of singular values divided by the largest one, i.e., $||\mathbf{\sigma}||_{1}/||\mathbf{\sigma}||_{\infty}$.
A large IA indicates a balanced distribution of singular values and a robust embedding, while a small IA indicates an imbalanced distribution and dimensional collapse (Guo et al., 2023b; Chi et al., 2022).
@@ 2.2. Sufficient Parameter Learning
Figure 2. Information Abundance ($log(IA)$) dynamics of SDEM’s bottom layer of experts and item embeddings on Kuairand1k. High and Low denote the items with the highest and lowest frequencies.
Taking SDEM as an example, we analyzed its training IA dynamics of experts and embeddings in Fig. 2. In Kuairand1k, S1 and S6 are the domains with the most and least data. In Fig. 2(a), the S6-specific expert exhibits a lower IA compared to S1, indicating that domains with a smaller data volume are more susceptible to dimensional collapse at the expert level. In Fig. 2(b), we further categorized items into five groups based on their exposure frequency and observed the collapse of item ID embeddings for both the highest and lowest frequency groups. We discovered that domain S6 exhibits a lower IA than the overall S6, indicating a more severe collapse, regardless of the group. Within each domain, the low-exposure group consistently shows a lower IA than the high-exposure group. We conclude with the following finding:
Finding 2. Models with domain-specific embeddings suffer from insufficient learning, indicated by dimensional collapse on embeddings of small domains, due to imbalanced data volume across domains.
@@
3. Our Method: Crocodile
We propose a multi-embedding architecture with Cross-experts Covariance Loss for Disentangled Learning (Crocodile) to tackle the above-mentioned challenges. It consists of multiple embeddings and experts, with a disentanglement loss upon them, domain-specific towers, and a customized gating mechanism connecting the experts and the towers, as shown in Fig. 3.